# Condensation of Composite Bosonic Trions in Interacting Bose-Fermi Mixtures

### v3

@@ Abstract
The Fourier transforms of the correlation functions yield the momentum distributions $n_{q}(k)=\frac{1}{L}\sum_{i,j}G_{q}(i,j)e^{ik(i-j)}$ for different particles $q$. In the TLL phase, paired fermions exhibit quasi-condensation around momentum $k=0$ at $t_{ff}=-3.0$ (alternatively, around $k=\pi$ at $t_{ff}=+3.0$, as shown in SM ). Similarly, the momentum distribution of bosons displays a sharp peak $k=0$, as depicted in Fig. 4(a). On the contrary, in the composite trion phase, the momentum distributions of both paired fermions and bosons are significantly broadened. Notably, only the composite trions exhibit a prominent peak in momentum space, which is visible in Figure 4(b). The condensation momentum of composite trions is $\pi$ for positive $t_{ff}$ and 00 for negative $t_{ff}$, depending on the primary distribution of the constituent fermion pairs.
Influence of magnitude of $t_{ff}$.— At small $|t_{ff}/t_{f}|$, another type of two-component Luttinger liquid phase emerges. In this phase, all relevant correlation functions—$G_{f},G_{b},G_{bf},G_{ff}$ and $G_{ffb}$—exhibit power-law decay (see Supplemental Materials SM ). This behavior contrasts with the case of small $|U_{bf}|$ and large $|t_{ff}|$, where $G_{f}$ and $G_{bf}$ decay exponentially, while $G_{ff}$, $G_{b}$, and $G_{ffb}$ show algebraic decay. When $|t_{ff}/t_{f}|>1.6$, the composite trion phase emerges, with only composite bosonic trions displaying QLRO. This indicates the importance of preformed fermion pairs in the formation of composite bosonic trions, despite the absence of QLRO for paired fermions within the composite trion phase.
Extended Bose-Fermi Hubbard Model.— At this point, our analysis based on the Bose–Fermi pair–hopping model has underscored the pivotal role of preformed fermion pairs in driving the composite bosonic trion phase. Motivated by these insights, we now extend our investigation to a more general density-density interaction framework. In particular, we investigate the extended Bose–Fermi Hubbard model Baranov et al. (2012); Wang et al. (2010); Orignac et al. (2010); Gómez-Lozada et al. (2025), which describes a BFM of dipolar particles interacting via both short-range and long-range dipole–dipole interactions. This model incorporates both first-neighbor ($V_{f1}$) and second-neighbor ($V_{f2}$) density-density interactions between fermions. Moreover, bosons and fermions interact via onsite ($U_{bf}$) and nearest-neighbor density-density interactions ($V_{bf}$). The extended Bose–Fermi Hubbard Hamiltonian reads
@@
This extended model not only captures the essential physics revealed by the analysis based on the Bose–Fermi pair–hopping model (1), but also provides a versatile platform for exploring the stability and robustness of the composite bosonic trion phase.
@@
The fermionic sector of the Hamiltonian (4) is a paradigmatic setup in one-dimensional lattice models of spinless fermions Mattioli et al. (2013); Dalmonte et al. (2015); Kane et al. (2017); Hohenadler et al. (2012); He et al. (2019); Gotta et al. (2021c). Its ground state supports a cluster Luttinger liquid of paired fermions with gapped single-fermion excitations when $V_{f1}<0$ and $V_{f2}>0$ at fermion densities below half-filling. We set $t_{f}=1,V_{f1}=-V_{f2}=-5.0$ and $\rho_{f}=1/5$, a typical parameter choice in the pair cluster Luttinger liquid, to facilitate the formation of preformed fermion pairs. The attraction between bosons and fermions can drives the system into the composite trion phase. As shown in Fig. 5, the trion correlation function $G_{ffb}$ exhibits QLRO with power-law decay, while the trion occupation number $n_{ffb}(k)$ features a pronounced peak at zero momentum. As the boson-fermion attraction strengthens, the system transitions from the TLL phase to the composite trion phase, as reflected in the central charge behavior in Fig. 5(c). The trion condensate remains stable over a broad range of interaction strengths, specifically for $-5.8\leq U_{bf}\leq-1.2$. Beyond this regime, the system becomes unstable and undergoes phase separation (PS), where clustered particles are randomly distributed without a well-defined central charge. Additionally, the trion condensate is favored when $t_{b}<t_{f}$, corresponding to heavy bosons and light fermions. For a fixed $U_{bf}=-V_{bf}=-5.0$, the composite trion phase remains robust within $0.4\leq t_{b}/t_{f}\leq 4.4$ SM .
@@
Discussions.— Our study unravels a general mechanism for forming a condensate of composite bosonic trions. Rather than relying on complex three-body interactions, our approach leverages the combination of preformed fermion pairs with bosons via simple two-body boson–fermion interactions. For the Bose–Fermi pair–hopping model, recent advances in Floquet engineering enable the simulation of pair hopping by enhancing pair tunneling while simultaneously suppressing single-particle tunneling Klemmer et al. (2024); Eckardt (2017); Fölling et al. (2007). Furthermore, Feshbach resonances provide a convenient means to tune boson–fermion attractions Schfer et al. (2020); Gross and Bloch (2017); Bloch et al. (2008), making the implementation of the pair–hopping model experimentally promising. In parallel, the extended Bose–Fermi Hubbard model—describing dipolar bosons and fermions confined in a one-dimensional optical lattice—offers an alternative route for realizing composite trion condensates. Combined with Feshbach resonance techniques, both short-range onsite interactions and long-range dipole–dipole interactions (including first- and second-neighbor density–density couplings) can be precisely controlled Baranov et al. (2012); Schfer et al. (2020); Gross and Bloch (2017). Notably, the composite trion phase occupies a substantial portion of the phase diagrams for both models, further enhancing its experimental feasibility. The formation of composite trions can be identified by the radio-frequency (rf) dissociation spectrum, and their binding energy can be measured based on the rf loss spectrum Yang et al. (2022a, b); Cao et al. (2024). Furthermore, the condensation of bosonic trions is expected to manifest in measurements of the pair momentum distribution Ries et al. (2015); Kuklov and Moritz (2007); Orso et al. (2010). Meanwhile, the suppression of superconducting correlations of fermions can be detected using interferometric techniques Carusotto and Castin (2005).
@@
In conclusion, we have demonstrated that a quasi–long–range coherent ground state of composite bosonic trions, characterized by a central charge $c=1$, emerges robustly in both the Bose–Fermi pair–hopping model and the extended Bose–Fermi Hubbard model. This phase is stabilized by a combination of negative binding energy, gapped single–particle excitations, and suppressed correlations of paired fermions, which together underpin the quasi–condensation of the bosonic trions and illuminate the underlying pairing mechanism between bosons and fermions. Our results also confirm the existence of the trion phase in multi-leg ladder systems beyond the 1D chain geometry SM . While our calculations have been conducted for simplicity on spinless fermions and bosons, we believe our results can be generalized to spinful systems, resulting in a richer phase diagram. Overall, our findings provide new insights into the three-body pairing mechanism among particles with different statistics and paves avenues for future experimental and theoretical investigations into intriguing pairing phases.
@@
Acknowledgements.— We thank Ting-Kuo Lee and Kai Chang for their invaluable suggestions. This work is supported by the National Key Research and Development Program of China Grant No. 2022YFA1404204, and the National Natural Science Foundation of China Grant No. 12274086.

### v4

@@ II.3 Influence of magnitude of $t_{ff}$
III Composite bosonic trions in the Extended Bose-Fermi Hubbard Model
At this point, our analysis based on the pair–hopping model has underscored the pivotal role of preformed fermion pairs in driving the composite bosonic trion phase. Motivated by these insights, we now extend our investigation to a more general density-density interaction framework. In particular, we investigate the extended Bose–Fermi Hubbard model Baranov et al. (2012); Wang et al. (2010); Orignac et al. (2010); Gómez-Lozada et al. (2025), which describes a BFM of dipolar particles interacting via both short-range and long-range dipole–dipole interactions. This model incorporates both first-neighbor ($V_{f1}$) and second-neighbor ($V_{f2}$) density-density interactions between fermions. Moreover, bosons and fermions interact via onsite ($U_{bf}$) and nearest-neighbor ($V_{bf}$) density-density interactions. The extended Bose–Fermi Hubbard Hamiltonian reads
@@ III Composite bosonic trions in the Extended Bose-Fermi Hubbard Model
This extended model not only captures the essential physics revealed by the analysis based on the pair–hopping model (1), but also provides a versatile platform for exploring the stability and robustness of the composite bosonic trion phase.
@@
The fermionic sector of the Hamiltonian (5) is a paradigmatic setup in one-dimensional lattice models of spinless fermions Mattioli et al. (2013); Dalmonte et al. (2015); Kane et al. (2017); Hohenadler et al. (2012); He et al. (2019); Gotta et al. (2021c). Its ground state supports a cluster Luttinger liquid of paired fermions with gapped single-fermion excitations when $V_{f1}<0$ and $V_{f2}>0$ at fermion densities below half-filling. We set $t_{f}=1,V_{f1}=-V_{f2}=-5.0$ and $\rho_{f}=1/5$, a typical parameter choice in the pair cluster Luttinger liquid, to facilitate the formation of preformed fermion pairs. The attraction between bosons and fermions can drive the system into the composite trion phase. As shown in Fig. 6, the trion correlation function $G_{ffb}$ exhibits QLRO with power-law decay, while the trion occupation number $n_{ffb}(k)$ features a pronounced peak at zero momentum. As the boson-fermion attraction strengthens, the system transitions from the TLL phase to the composite trion phase, as reflected in the central charge behavior in Fig. 6(c). The trion condensate remains stable over a broad range of interaction strengths, specifically for $-5.8\leq U_{bf}\leq-1.2$. Beyond this regime, the system becomes unstable and undergoes phase separation (PS), where clustered particles are randomly distributed without a well-defined central charge. Additionally, the trion condensate is favored when $t_{b}<t_{f}$, corresponding to heavy bosons and light fermions. For a fixed $U_{bf}=-V_{bf}=-5.0$, the composite trion phase remains robust within $0.4\leq t_{b}/t_{f}\leq 4.4$, as shown in Fig. 7.
Figure 7: Central charge $c$ as a function of $t_{b}/t_{f}$ in an $L=70$ system with PBC. Within the composite trion phase, $c=1$ for $0.4\leq t_{b}/t_{f}\leq 4.4$. ”PS” denotes phase separation for $t_{b}/t_{f}<0.4$, characterized by randomly distributed particle clusters lacking a well-defined central charge. The parameters are set as $V_{f1}=-V_{f2}=-5.0,U_{bf}=-V_{bf}=-5.0,U_{bb}=8.0$, and $\rho_{f}=2\rho_{b}=1/5,t_{f}=1$.
IV Discussions
@@
Our study unravels a general mechanism for forming a condensate of composite bosonic trions. Rather than relying on complex three-body interactions, our approach leverages the combination of preformed fermion pairs with bosons via simple two-body boson–fermion interactions. For the pair–hopping model, recent advances in Floquet engineering enable the simulation of pair hopping by enhancing pair tunneling while simultaneously suppressing single-particle tunneling Klemmer et al. (2024); Eckardt (2017); Fölling et al. (2007). Furthermore, Feshbach resonances provide a convenient means to tune boson–fermion attractions Schfer et al. (2020); Gross and Bloch (2017); Bloch et al. (2008), making the implementation of the pair–hopping model experimentally promising. In parallel, the extended Bose–Fermi Hubbard model—describing dipolar bosons and fermions confined in a one-dimensional optical lattice—offers an alternative route for realizing composite trion condensates. Combined with Feshbach resonance techniques, both short-range onsite interactions and long-range dipole–dipole interactions (including first- and second-neighbor density–density couplings) can be precisely controlled Baranov et al. (2012); Schfer et al. (2020); Gross and Bloch (2017). Notably, the composite trion phase occupies a substantial portion of the phase diagrams for both models, further enhancing its experimental feasibility. The formation of composite trions can be identified by the radio-frequency (rf) dissociation spectrum, and their binding energy can be measured based on the rf loss spectrum Yang et al. (2022a, b); Cao et al. (2024). Furthermore, the condensation of bosonic trions is expected to manifest in measurements of the pair momentum distribution Ries et al. (2015); Kuklov and Moritz (2007); Orso et al. (2010). Meanwhile, the suppression of superconducting correlations of fermions can be detected using interferometric techniques Carusotto and Castin (2005).
@@
In conclusion, we have demonstrated that a quasi–long–range coherent ground state of composite bosonic trions, characterized by a central charge $c=1$, emerges robustly in both the pair–hopping model and the extended Bose–Fermi Hubbard model. This phase is stabilized by a combination of negative binding energy, gapped single–particle excitations, and suppressed correlations of paired fermions, which together underpin the quasi–condensation of the bosonic trions and illuminate the underlying pairing mechanism between bosons and fermions. Our results also confirm the existence of the trion phase in multi-leg ladder systems beyond the 1D chain geometry (see Appendix C). While our calculations have been conducted for simplicity on spinless fermions and bosons, we believe our results can be generalized to spinful systems, resulting in a richer phase diagram. Overall, our findings shed light on the three-body pairing mechanism in Bose-Fermi mixtures and open avenues for future experimental and theoretical investigations into intriguing pairing phases.
Acknowledgements
@@
We thank Ting-Kuo Lee and Kai Chang for their invaluable suggestions. This work is supported by the National Key Research and Development Program of China Grant No. 2022YFA1404204, and the National Natural Science Foundation of China Grant No. 12274086.
Appendix A Detailed numerical setup
The numerical calculations employed the state-of-the-art Density Matrix Renormalization Group (DMRG) method, as referenced in White (1992); Fishman et al. (2022). Our simulations studied system sizes of up to $L=160$ with open boundary conditions (OBC) and $L=96$ with periodic boundary conditions (PBC). The computations involved approximately 150 sweeps, retaining $m=1500$ states for OBC with a typical truncation error $\epsilon\sim 10^{-9}$, and $m=3000$ states for PBC with $\epsilon\sim 10^{-8}$. Due to the substantial onsite repulsion ($U_{bb}=8.0$) and the relatively low boson density ($\rho_{b}=N_{b}/L=1/8$), it was adequate to set the maximum boson occupation number per site ($n^{cut}_{b}$) at 2. Notably, results for $n^{cut}_{b}\geq 1$ were comparable.
Appendix B Results for the Bose–Fermi model incorporating fermion pair hopping
Figure S1: Schematic phase diagram for the Bose–Fermi model incorporating fermion pair hopping in the $t_{ff}$-$U_{bf}$ parameter space for $\rho_{f}=2\rho_{b}=1/4$. The composite trion and TLL phases are the same as those described in the main text. ”TLL” denotes a two-component Luttinger liquid of bosons and paired fermions, whereas ”TLLBF” denotes a two-component Luttinger liquid of bosons and unpaired fermions. Solid red and blue circles represent fermions and bosons, respectively. Particles enclosed within black circles and ellipses indicate paired configurations. Negative values of $t_{ff}$ yield similar results to those for positive $t_{ff}$.
Figure S2: Central charge $c$ as a function of $t_{b}/t_{f}$ in the $L=160$ system with OBC. Within the composite trion phase, $c=1$ for $0.5\leq t_{b}/t_{f}\leq 4.8$. ”PS” denotes phase separation for $t_{b}/t_{f}<0.5$, characterized by randomly distributed particle clusters lacking a well-defined central charge. The parameters are set as $t_{f}=1,t_{ff}=3.0,U_{bf}=-2.8,U_{bb}=8.0$, and $\rho_{f}=2\rho_{b}=1/4$.
1 Schematic phase diagram as a function of $t_{ff}$ and $U_{bf}$
We present a qualitative ground-state phase diagram as a function of $t_{ff}$ and $U_{bf}$ for fermion and boson densities $\rho_{f}=2\rho_{b}=1/4$, as shown in Fig. S1. Negative values of $t_{ff}$ yield results similar to those for positive $t_{ff}$.
For small fermion pair-hopping amplitudes $t_{ff}$, the system approaches the limit of the standard Bose–Fermi Hubbard model Pollet et al. (2006). In this regime, the ground state is a two-component Luttinger liquid of bosons and unbound fermions, exhibiting a central charge of $c=2$. We refer to this as the “TLLBF” phase, where the single-particle correlation functions of both fermions and bosons display dominant quasi-long-range order. Due to the fermion density being twice that of the bosons, the TLLBF phase persists even at large boson–fermion interactions $U_{bf}$. In contrast, when the densities of the two species are equal and $U_{bf}$ is large, the ground state becomes a single-component Luttinger liquid of composite fermions formed by binding one fermion with one boson Pollet et al. (2006). Excessively strong $U_{bf}$ at $\rho_{f}=2\rho_{b}$ drives the system into a phase-separated state, where particle clusters distribute in distinct regions of the real-space lattice. This behavior can be intuitively understood as follows: at $\rho_{f}=2\rho_{b}$, unbound fermions and composite fermions coexist in the system, and the strong scattering of composite fermions by the remaining unbound fermions leads to phase separationSong et al. (2024). This regime lies beyond the scope of the schematic phase diagram. For instance, phase separation occurs at $U_{bf}=-28$ for $t_{ff}=-0.2$ with parameters $t_{f}=t_{b}=1$ and $U_{bb}=8$.
Figure S3: Momentum distribution functions $n(k)$ for $t_{ff}=+3.0,\rho_{f}=2\rho_{b}=1/4$ in a system of length $L=80$ with PBC. Panel (a) displays the results for the TLL phase at $U_{bf}=-0.3$, while panel (b) shows the results for the composite trion phase at $U_{bf}=-6.0$. The grey dashed lines guide the eyes at $k=\pi$.
For large $t_{ff}$ and small attractive $U_{bf}$, the system resides in a two-component Luttinger liquid phase consisting of bosons and paired fermions, denoted as the TLL phase in the main text. This TLL phase also appears in the repulsive $U_{bf}$ regime. When $U_{bf}=0$ and $t_{ff}$ is positive, the system exhibits a coexistence phase comprising neighboring paired fermions in a sea of unpaired fermions Gotta et al. (2021a). However, as $U_{bf}$ increases, this coexistence region rapidly narrows and eventually disappears. Furthermore, since this phase has negligible influence on the composite trion phase and occupies only a very small parameter region, it is not shown in the schematic phase diagram.
Additionally, for large $t_{ff}$ and strongly attractive $U_{bf}$, the composite trion phase occupies a broad region of the phase diagram. In this phase, only the composite bosonic trions exhibit dominant quasi-long-range order.
2 Influence of magnitude of $t_{b}/t_{f}$
The appearance of the trion condensate is facilitated when $t_{b}<t_{f}$, but it also remains stable over a wide range of $t_{b}\geq t_{f}$. For example, when $U_{bf}=-2.8t_{f}$, we observe that for $0.5\leq t_{b}/t_{f}\leq 4.8$, the system consistently remains in the composite trion phase, as shown in Fig. S2. This demonstrates that the hopping amplitudes $t_{b}$ and $t_{f}$ do not require highly precise tuning, and the trion condensate can emerge over a broad range of values. For too small $t_{b}/t_{f}$, the system becomes unstable to phase separated phase (PS), where clustered particles are randomly distributed without a well-defined central charge.
Figure S4: Correlation functions for (a) the TLL phase at $U_{bf}=-0.3$ and (b) the composite trion phase at $U_{bf}=-3.0$ in a 1D system of length $L=160$ with OBC. The parameters $t_{ff}=-3.0,\rho_{f}=2\rho_{b}=1/4$ are fixed and $r$ represents the distance from the reference site at $L/4$. Both (a) and (b) use a double-logarithmic scale. The solid lines denote power-law fitting to $|G(r)|\sim r^{-\alpha_{q}}$, where $\alpha_{b}\simeq 0.55,\alpha_{ff}\simeq 0.71,\alpha_{ffb}\simeq 1.13$ in (a), and $\alpha_{fbf}\simeq 1.05$ in (b). (c) The central charge is obtained by fitting the von Neumann entropy as a function of $U_{bf}$. In the composite trion phase, $c=1$ when $U_{bf}\leq-1.4$.
Figure S5: Fitting of the central charge $c$ from the von Neumann entanglement entropy $S_{vN}$ in the $L=160$ system with OBC at $t_{ff}=3.0$. Here, $l$ represents the subsystem length. The open squares depict numerical data, while the solid lines are fitted using Eq. (6).
3 Results for Negative $t_{ff}=-3.0$
In the limit of paired fermions ($|t_{ff}|\gg t_{f}$), the sign of $t_{ff}$ primarily affects the momenta where fermion pairs quasi-condensate, specifically at $k=0$ for $t_{ff}<0$ and $k=\pi$ for $t_{ff}>0$). Regardless of the sign of $t_{ff}$, bosons distribute predominantly around $k=0$ in momentum space. Since paired fermions are constituent parts of composite trions, the condensation momenta of the composite trions are either $k=0$ for $t_{ff}<0$ (see Fig. 4 in the main text) or $k=\pi$ for $t_{ff}>0$ (see Fig. S3(b)). The evolution of the correlation functions and central charge for $t_{ff}=-3.0$ display similar behaviors as those for $t_{ff}=+3.0$, which can be seen in Fig. S4 here and Fig. 2 in the main text. For $t_{ff}=-3.0$, the system fully transmutes into the composite trion phase with a central charge equal to 1 when $U_{bf}\leq-1.4$, which slightly differs from the case for $t_{ff}=+3.0$.
Figure S6: Charge density profiles $n_{i}$ at site $i$ for $t_{ff}=3.0,\rho_{f}=2\rho_{b}=1/4$ at various $U_{bf}$ values in the $L=160$ system with OBC. ”f” and ”b” denote fermions and bosons, respectively.
4 Entanglement entropy and central charge
[…]
$S_{vN}(l)=$ $\frac{c}{6}ln(\frac{2L}{\pi}sin(\frac{\pi l}{L}))$ (6)
$+B(\langle c^{\dagger}_{l}c_{l+1}\rangle+H.c.)+S_{0},$
where $S_{0}$ and $B$ are model-dependent fitting parameters, and $L$ denotes the system size. Due to the presence of additional oscillations in $S_{vN}(l)$, the local fermion hopping terms $B(\langle c^{\dagger}_{l}c_{l+1}\rangle+H.c.)$ are incorporated, with $B$ depending on the fitting Gotta et al. (2021a, c); Roux et al. (2009). In the TLL phase, $c=2$ corresponds to the two gapless modes: the Luttinger liquid of paired fermions and that of bosons. In the composite trion phase, $c=1$ corresponds to a single bosonic gapless mode of the Luttinger liquid of composite trions. Fig. S5 shows several typical fits of the central charge $c$ from the von Neumann entanglement entropy $S_{vN}$.
Figure S7: Density-density correlations with $t_{ff}=3.0$ for (a) the TLL phase at $U_{bf}=-0.3$ and (b) the composite trion phase at $U_{bf}=-3.0$ in a 1D system of length $L=160$ with OBC. $\rho_{f}=2\rho_{b}=1/4$ are fixed, and $r$ represents the distance from the reference site at $L/4$. A double-logarithmic scale is used in both (a) and (b). Solid lines denote power-law fitting to $|G(r)|\sim r^{-\alpha_{q}}$ where $\alpha_{b}\simeq 2.28,\alpha_{f}\simeq 1.64$ in (a) and $\alpha_{f}\simeq 1.27,\alpha_{b}\simeq 1.27$ in (b).
Figure S8: Structure factors of density correlation functions for a periodic system with $L=80$ for (a) fermions and (b) bosons at various $U_{bf}$ values.
5 Charge density waves
In 1D chains with open boundary conditions, Friedel oscillations can be observed in the charge density profilesMoreno et al. (2011). A $2k_{F}\ (k_{F}=\pi\rho_{f})$ charge-density-wave (CDW) order persists in both the TLL phase and the composite trion phase, as shown in Fig. S6. The $2k_{F}$ periodicity of CDW, rather than $k_{F}$, arises due to bound fermion pairs at $|t_{ff}|=3.0$. Bosons exhibit the same oscillatory behavior as fermions due to the onsite attractive interaction $U_{bf}$. At small $|U_{bf}|$ values, the density-density correlation functions of both species exhibit algebraic decay with power exponents close to $2$, indicating weak CDW order (see Fig. S7). As $U_{bf}$ increases, the amplitude of CDW is enhanced and the density-density correlations decay slower with power exponents approaching $1$, indicating a quasi-long-range CDW order. The decay exponents of both species are nearly identical, suggesting that bosons are locked to fermions in the strong coupling regime.
To further characterize the composite trion phase, we calculate the charge-density correlation functions $D_{q}(r)=\langle n^{q}_{i}n^{q}_{j}\rangle-\langle n^{q}_{i}\rangle\langle n^{q}_{j}\rangle$ for $q=f,b$. Based on the bosonization prediction for 1D systemsGotta et al. (2021c), the lowest order contributions to $D_{q}(r)$ have the form
$D_{q}(r)=\frac{A_{1}}{r^{2}}+A_{2}\frac{cos(2\pi\rho r)}{r^{2K_{q}}},$ (7)
where $A_{1},A_{2}$ are model-dependent amplitudes, and $K_{q}$ is the Luttinger parameter. In both the TLL phase and the composite trion phase, fermions are always in a paired configuration, hence $\rho=\rho_{f}/2=\rho_{b}$ for $D_{f}(r)$. Upon Fourier transformation, the density structure factor $D_{f}(k)$ display a $2k_{c}$ cusp with $k_{c}=\pi\rho$, signaling a $2k_{c}$ charge-density wave. Additionally, for bosons’ $D_{b}(k)$, the $2k_{c}$ cusp is only present in the composite trion phase and absent in the TLL phase. Furthermore, in Fig. S8(b), the $2k_{c}$ cusp becomes more prominent as the absolute value of $U_{bf}$ increases, indicating an enhanced tendency to form composite trions.
6 Results for fermions at half filling: $\rho_{f}=2\rho_{b}=\frac{1}{2}$
Figure S9: The central charge for a system with $\rho_{f}=2\rho_{b}=1/2$ at $t_{ff}=3.0$ in a 1D system of length $L=160$ with OBC. In the composite trion phase, $c=1$ when $U_{bf}\leq-2.0$.
The results obtained at another typical particle density, $\rho_{f}=2\rho_{b}=\frac{1}{2}$, are similar to those at $\rho_{f}=2\rho_{b}=\frac{1}{4}$, except for a larger value of $U_{bf}$ where the system fully transitions into the composite trion phase with a central charge of 1. For $\rho_{f}=2\rho_{b}=\frac{1}{2}$, the central charge $c=1$ when $U_{bf}\leq-2.0$, whereas for $\rho_{f}=2\rho_{b}=\frac{1}{4}$, $c=1$ when $U_{bf}\leq-1.2$. The gradual changes in the central charge as $U_{bf}$ varies are illustrated in Fig. S9. The fermion-boson attraction $U_{bf}$ drives the crossover from the TLL phase with $c=2$ to the composite trion phase with $c=1$.
Appendix C Stability of composite trions in ladder systems
The composite trion phase is not confined to 1D chain systems but can also stabilize in multi-leg ladder configurations. The Hamiltonian is designed to allow individual fermions and bosons to hop in both the $\hat{x}-$ and $\hat{y}-$directions, while the paired fermion hopping terms are restricted to the $\hat{x}-$direction. The system adopts a cylinder geometry with periodic and open boundary conditions in the $\hat{y}$ and $\hat{x}$ directions. Bosons interact with fermions through a local density-density attraction $U_{bf}$. Correlation functions are averaged over the number of legs $L_{y}$ as $G_{O}(x)=\frac{1}{L_{y}}\sum^{L_{y}}_{y=1}\langle O^{\dagger}(x_{0},y)O(x_{0}+x,y)\rangle$ with the reference site at $x_{0}\sim L_{x}/4$. Fig. S10 demonstrates that, among various correlations, only the composite trion correlation functions decay the slowest, following a power-law manner.
Figure S10: Correlation functions in the composite trion phase at $t_{ff}=3.0,U_{bf}=-12.0,U_{bb}=8.0$ for (a) 2-leg ladder with $L_{x}=48$ and (b) 3-leg cylinder with $L_{x}=32$. $x$ is the distance from the reference site $L_{x}/4$ in the $\hat{x}-$direction. Double-logarithmic scale is used. The solid lines denote power-law fitting to $|G(r)|\sim r^{-\alpha_{ffb}}$.
Appendix D composite fermion pairing phase at $N_{f}=N_{b}$
Figure S11: Correlation functions with $N_{f}=N_{b}$ for (a) the TLL phase at $U_{bf}=-0.4$ and (b) the composite fermion pairing (CFP) phase at $U_{bf}=-4.0$ in a 1Dsystem of length $L=160$ with OBC. Here, $t_{ff}=3.0,\rho_{f}=\rho_{b}=1/4$ are fixed, and $r$ represents the distance from the reference site at $L/4$. A double-logarithmic scale is used in both (a) and (b). The solid lines epresent power-law fitting to $|G(r)|\sim r^{-\alpha_{q}}$ where $\alpha_{b}\simeq 0.44,\alpha_{ff}\simeq 0.76,\alpha_{ffb}\simeq 1.13$, $\alpha_{bfbf}\simeq 2.4$ in (a) and $\alpha_{bfbf}\simeq 1.46$ in (b).
When the densities of fermions and bosons are equal, a composite fermion pairing phase emerges. A composite fermion consists of one fermion and one boson. At large $|U_{bf}|$ values, only the correlation function of paired composite fermions, given by
$G_{bfbf}(r)=\langle c^{\dagger}_{i}b^{\dagger}_{i}c^{\dagger}_{i+1}b^{\dagger}_{i+1}c_{j}b_{j}c_{j+1}b_{j+1}\rangle,r=|i-j|,$ (8)
demonstrates power-law decay behavior with exponents $\alpha_{bfbf}<2$, as shown in Fig. S11. Consequently, the dominant QLRO is characterized by the pairing of composite fermions. The pairing phase of composite fermions, consisting of one fermion and one bosonic hole, has been discussed earlier in the context of polaronsDanshita and Mathey (2013); Mathey et al. (2004); Mathey and Wang (2007b). Here, polarons refer to atoms of one species dressed by screening clouds formed by atoms of another species. The identification of this composite-fermion-paired superconducting phase further enriches the landscape of unconventional many-body pairing states in Bose–Fermi systems.